# Modulation of Non-equilibrium Structures of Active Dipolar Particles by an External Field

### v1

@@ I Introduction
In the context of active dipolar particles in external fields, experimental studies on suspensions of magnetotactic bacteria highlighted a rich diversity of collective behaviors, including the formation of tunable three-dimensional (3D) clusters [33], band- and vortex-like structures [34, 35, 36], magneto-convection patterns [37], and pearling instabilities [38]. Furthermore, suspensions of magnetic microswimmers with hydrodynamic interactions in external alignment fields have been studied using kinetic theory [21, 22], revealing emergent dynamical patterns such as sheets and columns traveling along the field direction, corroborating experimental observations. The two-dimensional (2D) self-assembly of active dipolar particles under an external alignment field has also been examined at various densities using Brownian dynamics simulations [23]. The analysis showed that, while weak fields have little to no influence on the systems, strong fields prevent the formation of interconnected structures at large dipolar interactions, favoring instead ordered columnar clusters aligned with the field [23]. For systems at intermediate density, the external field was found to strongly mitigate the effect of activity, in striking contrast with the zero-field case [12]. However, how the interplay between self-propulsion, anisotropic long-range dipolar interactions, and external alignment fields impact the collective structure formation in 3D systems remains completely unexplored. It would be interesting to see how activity and external orientational constraints affect the structure formation of dipolar particles compared to their passive counterparts [39, 40, 41].
@@ II.2 Units and simulation parameters
[…]
Brownian dynamics simulations were carried out using the LAMMPS simulation package [45] to implement and integrate equations (5) and (6), using a time-step of $dt^{*}=2\times 10^{-4}$. For each investigated value of dipolar coupling strength $\lambda$, the system was first equilibrated in the passive zero-field limit, i.e., $f^{a*}$= 0 and $\eta=0$, where the initial configuration was that of randomly oriented particles located on a cubic lattice. Active systems were then equilibrated by restarting simulations from the equilibrated configuration of passive systems, setting $f^{a*}$ to the probed values while keeping $\eta=0$. Finally, a homogeneous alignment field was applied along the $z$-direction of the equilibrated passive and active boxes by instantaneously varying $\eta$ from 0 to the investigated values. During equilibration runs, convergence of the total potential energy, as well as polarization when $\eta\neq 0$ were monitored to ensure the systems were equilibrated. Reaching a steady state typically took from $10^{2}\tau$ to $10^{5}\,\tau$, depending on the $\{\lambda,f^{a*},\eta\}$ combination. Production runs were subsequently performed on the equilibrated systems to compute the relevant structural and dynamical properties, with typical length of $10^{3}\tau$, unless otherwise stated.
II.3 Analysis tools
@@
II.3.1 Clustering analysis
@@
It is well-established that dipolar particles in 3D can self-assemble into chains, rings, and percolated network structures in the low-density regime [30, 46, 47, 48, 13]. To quantify the effect of an external field on self-assembly, we perform a clustering analysis of the configurations based on a simple distance criterion, similar to that used in prior studies of active dipolar particles [13, 12], using k-d trees [49] for an efficient nearest neighbor search. Specifically, two particles are considered bound if their center-to-center distance $r_{ij}$ is smaller than a threshold cutoff, $r_{b}$. The distance criterion was set at $r_{b}^{*}$= 1.2, based on the first minimum of the radial distribution function, almost insensitive to variations in $\lambda$, $f^{a*}$, or $\eta$. Identified clusters are then sorted according to three topologies: chains are clusters with more than two particles, in which there exist exactly two particles with one neighbor and all the other particles have two neighbors; rings are clusters in which every particle has exactly two neighbors; branched structures are clusters in which at least one particle has more than two neighbors.
@@
where $N_{\alpha}(n)$ is the number of clusters of type $\alpha$ with length $n$ particles, delimited-⟨$\langle$..$\rangle$ … ⟩ refers to an average over relevant configurations, and the minimum length $n_{\alpha}^{\text{min}}$ equals 3 for topology $c$ and $r$, 4 for topology $b$, and 1 for topology $s$. From equation (7) we can calculate the average size of clusters with topology $\alpha$ as
@@
II.3.2 Angular distribution and mean polarization
@@
II.3.3 Classification of steady-states
To distinguish between the different steady-states observed in simulations, we employ a classification based on the order parameters obtained from the clustering analysis [13]. More precisely, the term gas is employed to refer to systems where most particles are in a monomeric state, with limit fixed at $\phi_{p}<0.5$, further implying that $\phi_{\text{max}}<0.5$. The term string fluid is used to refer to systems in which more than $50\%$ of the particles are in polymerized form ($\phi_{p}\geq 0.5$) but without percolation, the threshold being set at $\phi_{\text{max}}=0.7$, aligning with visual inspections of the simulation boxes and previous studies [48, 13]. Lastly, a (system-spanning) percolated network is identified whenever $\phi_{\text{max}}\geq 0.7$, which further implies $\phi_{p}\geq 0.5$. To further differentiate between orientationally ordered and disordered states, we introduce the term “Polarized” as a prefix for the latter denominations whenever $P\geq 0.5$. The classification of structural states based on order parameters $\phi_{p}$, $\phi_{\text{max}}$, and $P$ is summarized in Tab. 1.
II.3.4 Dynamical features
In order to quantify the bond dynamics, we compute a bond time autocorrelation function (TACF) which measures the probability that a bond existing at time $t^{\prime}$ still exists at time $t^{\prime}+t>t^{\prime}$, quantifying how fast the structures rearrange. This is done by assigning a bond variable $n_{ij}(t)$ at each time $t$ for every pair of particles $i$ and $j$, as described in [39]. This variable is set to 1 if particles $i$ and $j$ are neighbors according to the distance criterion introduced in Sec. II.3.1, and 0 otherwise. The bond TACF is then calculated as

### v2

@@ I Introduction
In the context of active dipolar particles in external fields, experimental studies on suspensions of magnetotactic bacteria [33, 34] highlighted a rich diversity of collective behaviors, including the formation of tunable three-dimensional (3D) clusters [35], band- and vortex-like structures [36, 37, 38], magneto-convection patterns [39], and pearling instabilities [40]. Furthermore, suspensions of magnetic microswimmers with hydrodynamic interactions in external alignment fields have been studied using kinetic theory [21, 22], revealing emergent dynamical patterns such as sheets and columns traveling along the field direction, corroborating experimental observations. The two-dimensional (2D) self-assembly of active dipolar particles under an external alignment field has also been examined at various densities using Brownian dynamics simulations [23]. The analysis showed that, while weak fields have little to no influence on the systems, strong fields prevent the formation of interconnected structures at large dipolar interactions, favoring instead ordered columnar clusters aligned with the field [23]. For systems at intermediate density, the external field was found to strongly mitigate the effect of activity, in striking contrast with the zero-field case [12]. However, how the interplay between self-propulsion, anisotropic long-range dipolar interactions, and external alignment fields impact the collective structure formation in 3D systems remains completely unexplored. It would be interesting to see how activity and external orientational constraints affect the structure formation of dipolar particles compared to their passive counterparts [41, 42, 43].
@@ II.2 Units and simulation parameters
Simulations were performed on a system of N=10 648N=10\,648 dipolar ABPs in a cubic box of edge size $L^{*}$ with periodic boundary conditions (PBC) in all three spatial directions. We made sure to avoid significant interactions between periodic image by checking that the maximum magnitude of the dipolar potential $2\lambda/{r^{*}_{ij}}^{3}$ is much smaller than $10^{-3}$ at half-box length. A cut-off radius of $L^{*}/2$ was accordingly employed in the potential of Eq. (3) to ensure an efficient treatment of dipolar interactions. The pairwise steric interaction strength was set to $\epsilon^{*}$= 1, and the particle density to $\rho^{*}$= 0.02 (packing fraction $\Phi\approx 0.01$) to match the dilute conditions of Ref. [13]. The dipolar coupling strength $\lambda$, the active force $f^{a*}$ and the field coupling strength $\eta$ were varied in the ranges $1\leq\lambda\leq 12.25$, $0\leq f^{a*}\leq 300$, and $0\leq\eta\leq 10$, respectively.
Brownian dynamics simulations were carried out using the LAMMPS simulation package [47] to implement and integrate equations (5) and (6), using a time-step of $dt^{*}=2\times 10^{-4}$. For each investigated value of dipolar coupling strength $\lambda$, the system was first equilibrated in the passive zero-field limit, i.e., $f^{a*}$= 0 and $\eta=0$, where the initial configuration was that of randomly oriented particles located on a cubic lattice. Active systems were then equilibrated by restarting simulations from the equilibrated configuration of passive systems, setting $f^{a*}$ to the probed values while keeping $\eta=0$. Finally, a homogeneous alignment field was applied along the $z$-direction of the equilibrated passive and active boxes by instantaneously varying $\eta$ from 0 to the investigated values. During equilibration runs, convergence of the total potential energy, as well as polarization when $\eta\neq 0$ were monitored to ensure the systems were equilibrated. Reaching a steady state typically took from $10^{2}\tau$ to $10^{5}\,\tau$, depending on the $\{\lambda,f^{a*},\eta\}$ combination. Production runs were subsequently performed on the equilibrated systems to compute the relevant structural and dynamical properties, with typical length of $10^{3}\tau$, unless otherwise stated. We verified that the WCA excluded-volume interactions remain sufficiently hard-core even at large active forces, preventing any significant interparticle overlap, as confirmed by the radial pair distribution function (see Appendix B).
III Impact of external field on passive self-assembly
@@
(a)(b)
[…]
We conclude this section by presenting the phase-diagram of passive dipolar particles as a function of $\lambda$ and $\eta$, based on the following systematic classification. To do so, we employ a combination of three order parameters [12, 13]: the fraction of polymerized particles $\phi_{p}$, the fraction of particles in the largest cluster $\phi_{\text{max}}$, and the polarization $P$ (see Appendix A for definitions). More precisely, the term gas is employed to refer to systems where most particles are in a monomeric state, with limit fixed at $\phi_{p}<0.5$, further implying that $\phi_{\text{max}}<0.5$. The term string fluid is used to refer to systems in which more than $50\%$ of the particles are in polymerized form ($\phi_{p}\geq 0.5$) but without large-scale connectivity, the threshold being set at $\phi_{\text{max}}=0.7$, aligning with visual inspections of the simulation boxes and previous studies [50, 13]. Lastly, a (system-spanning) percolated network is identified whenever $\phi_{\text{max}}\geq 0.7$, further implying $\phi_{p}\geq 0.5$. To account for the effect of the external field, we differentiate between orientationally ordered and disordered states by introducing the term “Polarized” as a prefix for the latter denominations whenever $P\geq 0.5$. The classification of structural states based on order parameters $\phi_{p}$, $\phi_{\text{max}}$, and $P$ is summarized in Tab. 1.
In the limit of no activity, the field primarily induces a polarization of the pre-existing structures, with the system developing orientational order along the field direction. Thus, gases transform into polarized gases; string fluids transform into polarized string fluids; percolated networks first progressively polarize while retaining their interconnected configurations, and ultimately transform into polarized string fluids at stronger fields, indicating the loss of network structure in favor of independent columnar clusters, as observed in Fig. 2. Polarization naturally occurs earlier in systems with higher $\lambda$ values (with the exception of $\lambda=12.25$, as discussed above), owing to cooperative dipolar interactions. Furthermore, we find that the zero-field gas present at $\lambda=4$ transitions into a string fluid, and eventually into a polarized string fluid, with increasing $\eta$. This attests to the propensity of an external field to favor polymerization, as previously noted.
Figure 6: Steady-states diagrams of 3D dipolar ABPs at $\rho^{*}$= 0.02 as a function $\eta$ and $f^{a*}$ for dipolar coupling strengths $\lambda=$ (a) 4, (b) 6.25, (c) 9, and (d) 12.25. Symbols indicate different structural states based on Tab. 1 and are coded as in Fig. 4. The color bar shows the normalized polarization $P$ [Eq. (12)]. Dashed lines show the $P\geq 0.5$ limit. The x- and y-axes are shared across columns and rows, respectively.
(a)(b)(c)(d)
Upon turning on the external field (middle and top panels), interesting structures emerge when $\lambda$ is sufficiently large. At the intermediate value of $\lambda=6.25$, monomeric and short active polymeric structures present at zero-field self-assemble into longer chained topologies overall aligned with the direction of the field, which keep growing in size with further increasing $\eta$. At $\lambda=9$, the long chains and branched structures found at $\eta=0$ arrange into several elongated, branched columnar clusters spanning the whole simulation box, similar to the passive case. At the highest dipolar coupling strength of $\lambda=12.25$, the percolated network structure, observed in the zero-field limit, persists at intermediate field values $\eta\sim 5$, although it appears with branches aligned with the external field, before it eventually transforms into columnar structures at $\eta=10$. The effect of activity on the spatial ordering of columnar clusters is discussed in Appendix B.
Figs. 6(a)–(d) present the steady-state diagrams of dipolar ABPs as a function of $\eta$ and $f^{a*}$ for dipolar coupling strengths $\lambda=4$, 6.25, 9, and 12.25. The classification of states follows the criteria introduced in Sec. III. At the dipolar coupling value of $\lambda=4$, zero-field active systems remain in the gas state. As $\eta$ increases, field-induced polymerization is generally not observed, and systems simply transition to polarized gases. An exception is observed for $f^{a*}$= 10 where a transition toward a polarized string fluid occurs, albeit requiring a much higher field strength than in the passive case. Furthermore, the polarization, indicated by the color overlay, appears largely independent of activity. This contrasts with higher values of the dipolar coupling ($\lambda\geq 6.25$), where the $P\geq 0.5$ threshold lines (dashed lines) exhibit a dip at intermediate activity levels, indicating that the transition to polarized states shifts to higher field strengths compared to both low and high activity. Interestingly, this appears at values of $f^{a*}$ near the string fluid to gas transition, i.e., where the systems consists of many independent active polymerized structures. As will be discussed in Sec. IV.2, this change actually corresponds to a transition from an enhanced to a hindered polarization response.
IV.2 Field-induced polarization
Figure 7: Mean normalized polarization $P$ [Eq. (12)], as a function of the field coupling strength $\eta$ at dipolar coupling strength $\lambda=$ (a) 1, (b) 6.25, (c) 9, and (d) 12.25, and for active forces $f^{a*}$= 0, 10, 50, 100, and 300. Dashed lines correspond to the Langevin function $P_{L}$ [Eq. (14)]. The x- and y-axes are shared across columns and rows, respectively.
(a)(b)(c)(d)
Figure 8: Angular probability distribution $\psi(\theta)$ at ($\lambda$, $\eta$) = (a) ($6.25$, $0.7$), (b) ($6.25$, 55), (c) ($12.25$, $0.7$), (d) ($12.25$, 55), and active forces $f^{a*}$= 0, 10, 30, 80, and 300. Dashed lines correspond to the Langevin distribution $\psi_{L}$ [Eq. (13)]. The x-axis is shared across columns.
(a)(b)(c)(d)
Figure 9: Schematic of the polarization-loss mechanism of self-assembled dipolar ABPs (a) by reorientation, and (b) by bond-breaking.
(a)(b)
(a)(b)
We argue that the non-monotonic behavior observed in Figs. 7 and 8 arises from the existence of active short filamentous assemblies. As the active force increases, large interconnected branched structures break into smaller assemblies, e.g., active chains, that frequently reorient and restructure due to activity-induced bond breaking and reformation, causing the simultaneous depolarization of multiple dipoles. These structural correlations lead to greater polarization loss than in a gaseous state, resulting in sub-Langevin behavior. In Fig. 9, we present a schematic representation of these mechanisms in two dimensions. Further increasing activity results in the onset of a gas whose orientational responds corresponds to $P_{L}$, thereby giving rise to the observed non-monotonous behavior.
This interpretation is supported by Fig. 10, where we plot, in the top panels, the mean normalized polarization $P$, and in the bottom panels, the mean chain fraction $\phi_{c}$, as functions of $f^{a*}$ for $\lambda=$ (a) 6.25 and (b) 12.25, and for field strengths $\eta=0.7$, $5$, and $10$. In agreement with the above suggestion, we observe that the polarization minima coincide with the chain fraction maxima, emphasizing the role of structural correlations of chain-like elements in the polarization behavior. This maximum logically shifts to higher activities as $\eta$ increases and promotes bonding. However, although a maximum in $\phi_{c}$ is observed for both dipolar coupling strengths, the polarization minimum is considerably more pronounced at $\lambda=12.25$ than at $\lambda=6.25$.
IV.3 Orientational dynamics
To elucidate the non-monotonic polarization behavior highlighted in the previous section, we plot in Figs. 11(a)–(b) the orientational TACF $C_{e}(t)$ at field coupling strength $\eta=7$ and dipolar coupling strengths $\lambda=6.25$ and 12.25, for various active forces. After an initial short-time exponential decay (see insets in Fig. 11), the orientational TACFs exhibit a slower, non-exponential relaxation toward a plateau. By fitting $C_{e}(t)$ at short times $t\lesssim 0.015$ with an exponential function of the form $\exp[-2D_{r}^{\text{short}}t]$, we extract the short-time rotational diffusion coefficient $D_{r}^{\text{short}}$. Figs. 12(a)–(d) show the short-time rotational diffusion coefficient $D_{r}^{\text{short}}/D_{r}$ as a function of the active force $f^{a*}$ for field coupling strengths $\eta=0$, 2, 7, and 10, and various dipolar coupling strengths. For $\lambda=1$, $D_{r}^{\text{short}}/D_{r}$ remains independent of activity across all field strengths, consistent with the system being in a gas-like state. At higher dipolar coupling strengths, where self-assembled structures emerge, $D_{r}^{\text{short}}/D_{r}$ falls below unity at low $f^{a}$ due to orientational constraints imposed by particle bonding. As the active force increases and induces bond breaking, the gradual emergence of a gaseous phase restores the rotational diffusion characteristic of free particles. The long-time limit of $C_{e}(t)$, related to the mean polarization via $\lim_{t\to\infty}C_{e}(t)=|\langle\hat{\mathbf{e}}_{i}\rangle|^{2}$, varies non-monotonically with the active force $f^{a*}$. It first decreases to a minimum before approaching a saturation value, with the minimum more pronounced for $\lambda=12.25$ than for $\lambda=6.25$. This trend corroborates the behavior of the polarization discussed in Sec. IV.2, suggesting that these variations reflect changes in the structural state of the system.
Figure 11: Orientational time autocorrelation function $C_{e}(t)$ [Eq. (16)] at field coupling $\eta=7$ for active forces $f^{a*}$= 0, 10, 50, 100, and 300, and dipolar coupling strengths $\lambda=$ (a) 6.25 and (b) 12.25. The y-axis is in log-scale. Insets show the exponential decay of $C_{e}(t)$ at short-times.
(a)(b)
Figure 12: Short-time rotational diffusion coefficient $D_{r}^{\text{short}}$ as a function of the active force $f^{a*}$ at field coupling strengths $\eta=\textbf{(a)}~0$, (b) 2, (c) 7, and (d) 10, for dipolar coupling strengths $\lambda=1$, $6.25$, 99, and $12.25$. The x- and y-axes are shared across columns and rows, respectively.
(a)(b)(c)(d)
Given the highlighted interplay between structure and response to the external field, we examine in this section how the field influences structural changes. To this end, we plot in Figs. 13(a)–(d) the polymerization degree $\phi_{p}$ as a function of $\eta$ for dipolar coupling strengths $\lambda=$ 1, 4, 6.25, and 12.25, and various active forces. At the lowest value of $\lambda=1$, where the systems remain in a gas state, polymerization is largely independent of the field strength, showing only a slight increase with $f^{a*}$ due to enhanced two-body collisions [53, 9]. For higher values of the dipolar coupling strength $\lambda\geq 4$, polymerization increases significantly with $\eta$, provided that the systems are not fully polymerized in the zero-field limit and that the active force is sufficiently low compared to the strength of the dipolar forces and field strength. As previously mentioned in Sec. III, this mechanism arises from the alignment field stabilizing head-to-tail bonding configurations, in competition with activity, just as observed in the zero-field limit [12, 13].
(a)(b)(c)(d)
(a)(b)
(a)(b)
(a)(b)(c)
IV.5 Bond dynamics
Figure 17: Normalized bond TACF $C_{b}(t)/C_{b}(0)$ [Eq. 15], at $\lambda=12.25$ for field-strength $\eta=$ (a) 0, and (b) 10, and activities $f^{a*}$= 0, 10, 20, 30, 50, 100, and 200. Dashed lines connecting the markers represent stretched-exponential fits. The y-axis is shared by both subfigures.
(a)(b)
We find that the $C_{b}(t)$ curves can be well described by a stretched-exponential decay of the form $\exp[-(t/\tau_{K})^{\beta}]$, where $\tau_{b}=\tau_{K}\Gamma(\beta^{-1})/\beta$ can be interpreted as the mean bond lifetime [56], with $\Gamma$ the well-known gamma-function. The evolution of $\tau_{b}$ with activity at $\lambda=12.25$ is presented in Fig. 18 for field coupling strengths $\eta=0$ and $10$. Consistent with previous observations, increasing $f^{a*}$ leads to a substantial decrease in the mean bond lifetime. Conversely, $\tau_{b}$ values are systematically larger at $\eta=10$ than in the zero-field case, which corroborates the arguments made in Sec. IV.4 that the interplay between activity and the external field consists of a competition between two antagonistic effects for bond stability, with activity weakening bonds and the external field favoring their formation.
Figure 18: Mean bond lifetime $\tau_{b}$ as a function of the active force $f^{a*}$ at $\lambda=12.25$, for field coupling strengths $\eta=$ 0, and 10.
V Conclusions
Under an external field, we find that active assemblies exhibit a much richer phenomenology than their passive counterparts. In passive systems, the zero-field structures simply polarize along the direction of the field, leading to (i) an increased polymerization at low dipolar coupling, and (ii) the formation of independent columnar clusters at higher dipolar coupling, signifying the breakdown of the network structures. In contrast, in active systems, the orientational constraint imposed by the external field and self-propulsion can act cooperatively to favor connectivity. In particular, we find that (i) low activity levels extend the range of fields over which a percolated network persists, and (ii) at strong dipolar coupling strengths, a weak external field can induce system-spanning networks at intermediate activity levels. These results demonstrate the greater robustness of active networks to weak external fields compared to their passive counterparts, a consequence of continuous activity-induced dynamical rearrangements, where moderate orientational constraints enhance the likelihood of structural encounters during reorganization. Conversely, high fields disrupt network connectivity, leading to the formation of polarized string fluids. In addition, contrary to the passive case, active systems with strong dipolar coupling strengths exhibit a lower polarization than a system of independent dipoles, which we attribute to the presence of active self-assembled structures. In the large-activity regime, where binding is suppressed, the system forms a dipolar active gas with a polarization response following the Langevin function.
In the ongoing pursuit of strategies to control and program the self-assembly of active particles, our findings highlight the effectiveness of external fields in directing the structural and orientational organization of self-propelled dipolar systems. These results provide valuable guidelines for manipulating the collective behavior of magnetotactic bacteria [33, 34] via external fields. Moreover, recent experimental realizations of active dipolar particles in both two- and three-dimensions [57, 20, 58, 59], where self-propulsion arises from induced charge electrophoresis [60], simultaneously generating electric dipoles, exhibit closely related dynamics. We anticipate similar phenomenology for active dipolar colloids driven by electrophoresis. In such systems, however, the presence of ions screens the electrostatic potential, making the dipolar interactions effectively short-ranged. Moreover, one must consider the intrinsic coupling between dipole–dipole interactions and particle activity induced by the external field. Lastly, the interplay between hydrodynamic and dipolar interactions is expected to play an important role for dipolar microswimmers at appropriate concentration and dipolar strength regimes [61, 62, 22].
As a promising avenue for future research, exploring the hysteresis behavior of active percolated networks could open new possibilities for designing the next generation of smart, reconfigurable materials. In particular, the combination of activity and external fields may serve as an effective mechanism for memory reset, enabling controllable self-assembly and disassembly cycles. Another open question concerns the relationship between connectivity and rigidity percolation in active networks, and how it governs their emergent mechanical properties. More broadly, extending the present study to higher densities, where passive dipolar particles are known to spontaneously develop orientational order [30, 49], could reveal collective flocking states analogous to those observed in two-dimensional active dipolar systems [12].
Acknowledgements.
Appendix A Analysis details
A.1 Clustering analysis
@@
Dipolar particles in 3D can self-assemble into chains, rings, and percolated network structures in the low-density regime studied here [30, 48, 49, 50, 13]. To classify the states observed in simulations, we perform a clustering analysis of the configurations based on a simple distance criterion [13, 12]: two particles are considered bound if their center-to-center distance $r_{ij}$ is smaller than a threshold cutoff, $r_{b}$. The distance criterion was set at $r_{b}^{*}$= 1.2, based on the first minimum of the radial distribution function, almost insensitive to variations in $\lambda$, $f^{a*}$, or $\eta$. We used k-d trees [63] for an efficient nearest neighbor search. Identified clusters are then sorted according to three topologies: chains are clusters with more than two particles, in which there exist exactly two particles with one neighbor and all the other particles have two neighbors; rings are clusters in which every particle has exactly two neighbors; branched structures are clusters in which at least one particle has more than two neighbors.
@@
where $N_{\alpha}(n)$ is the number of clusters of type $\alpha$ with length $n$ particles, $\langle$$\rangle$ refers to an average over relevant configurations, and the minimum length $n_{\alpha}^{\text{min}}$ equals 3 for topology $c$ and $r$, 4 for topology $b$, and 1 for topology $s$. From equation (7) we can calculate the average size of clusters with topology $\alpha$ as
@@
A.2 Angular distribution and mean polarization
@@
A.3 Dynamical features
To quantify the bond dynamics, we compute a bond time autocorrelation function (TACF) which measures the probability that a bond existing at time $t^{\prime}$ still exists at time $t^{\prime}+t>t^{\prime}$, quantifying how fast the structures rearrange. This is done by assigning a bond variable $n_{ij}(t)$ at each time $t$ for every pair of particles $i$ and $j$, as described in [41]. This variable is set to 1 if particles $i$ and $j$ are neighbors according to the distance criterion introduced in Appendix A.1, and 0 otherwise. The bond TACF is then calculated as